OpenStudy (anonymous):

CAN SOMEONE HELP ME SIMPLIFY AND WRITE THIS INTO STANDARD FORM (4X^2+9X+1) + 2X^2+7X+13)

OpenStudy (anonymous):

combine like terms \[4x^2+2x^2=6x^2\] \[7x+9x=16x\] \[1+13=14\]

OpenStudy (anonymous):

answer is \[6x^2+16x+14\]

OpenStudy (anonymous):

I DID THAT BUT ON THE PAPER IT SAYS SIMPLIFY , I THOUGHT IT MEANING MULTIPLY

OpenStudy (anonymous):

if it meant multiply it would have been \[(4X^2+9X+1)(2X^2+7X+13)\]

OpenStudy (anonymous):

satellite73 ? can you help with something